OpenStudy (anonymous):

Ratios: An airline company oversells 6 tickets for ever 21 passengers. How many tickets will the airline oversell for 350 passengers? Ratios: An airline company oversells 6 tickets for ever 21 passengers. How many tickets will the airline oversell for 350 passengers? @Mathematics

OpenStudy (anonymous):

100 tickets

OpenStudy (anonymous):

6x350/21

OpenStudy (anonymous):

6/21=x/350 6x350/21=x 100=x

OpenStudy (anonymous):

100 tickets, they are in direct proportion.